# On the Limitations of Fine-tuned Judge Models for LLM Evaluation

### v2

@@ 3.5 No Benefits from Prompt Engineering
    In-context Learning (ICL): where task demonstrations are integrated into the prompt as illustration. In our work, we randomly select 2-4 ICL demonstrations from the training set based on the max-context length.
@@
    Chain-of-thought (CoT): where the input prompt is structured in a way that mimics human reasoning. In our work, the judge model is forced to generate the explanation first, then provide a final judgement.
3.6 The Essence of Fine-tuned Judge: A Task-specific Classifier
@@
Although prior research on instruction-tuning all emphasizes the importance of data diversity Zhou et al. (2023); Lu et al. (2024), the fine-tuning of LLM judges is doing the opposite thing. Therefore, after fine-tuning for a single task with a fixed prompt template, the model lost its generalization ability, and degenerate into a task-specific classifier, which exhibits several limitations due to overfitting.
4 When to Trust Fine-tuned Judges?
Model Method SALAD-Bench JudgeLM-test PandLM-test Auto-J-test Average
accuracy F1 accuracy F1 accuracy F1 agreement
JudgeLM-7B random 82.50 57.36 80.40 71.87 67.54 65.20 43.25 59.42
perplexity 88.09 59.57 85.48 74.38 72.34 63.25 49.57 61.69
SoftEnt 95.10 63.62 89.40 78.09 77.96 64.01 55.89 65.40
SoftEnt-Cali 91.56 61.47 88.70 76.30 79.76 67.05 56.32 65.29
PandaLM-7B random 56.67 36.67 68.09 48.26 66.73 58.86 39.37 45.79
perplexity 59.37 55.03 70.15 52.62 68.53 57.73 42.53 51.98
SoftEnt 63.13 62.07 72.96 53.67 75.55 62.53 43.53 55.45
SoftEnt-Cali 66.77 63.77 73.91 55.97 76.35 67.30 45.40 58.11
Auto-J-13B random 72.29 48.10 77.27 60.7 71.54 63.51 46.12 54.61
perplexity 77.47 52.56 80.28 61.85 75.55 65.84 50.35 57.65
SoftEnt 81.98 54.68 79.13 63.58 80.16 66.61 53.01 59.47
SoftEnt-Cali 82.08 54.75 80.19 63.58 81.36 70.00 52.30 60.16
Table 9: Comparison of different reliability indicators for the judge models. We split the test sets into halves based on different indicators, and report the performance of the judge on the half with higher scores. The on random-split baseline denotes that the results are averaged over three times.
Despite the limitations revealed in our study, we aim not to disregard the significance of fine-tuned judges entirely. While LLM exhibit excellent performance among various tasks, task-specific models are still wildly used everywhere in natural language processing. Therefore, it deserves our discussion about the reliability of the fine-tuned judges. For this purpose, we propose a quantitative indicator to estimate whether an sample can be reliably evaluated by a judge model.
Borrowing the idea from confidence estimation Huang et al. (2024), we propose to quantify the reliability of the judge based on softmax entropy. Given an instruction $x$ and a fine-tuned judge model with parameters $\theta$, the reliability of generating response $y$ can be factorized as:
$\textrm{SoftEnt}(y|x,\theta)=-\frac{1}{T}\sum_{t=1}^{T}\sum_{v=1}^{V}p_{\theta}(y_{t}^{v})\mathrm{log}p_{\theta}(y_{t}^{v})$
where $p(y_{t})$ represents the conditional distribution $p(y_{t}|x,y_{<t},\theta)$ at each decoding step, $T$ is the response length, and $V$ is the vocabulary size.
As we would like to quantify whether the sample lies in the task-specific fine-tuning scope, we further calibrate the SoftEnt as follows:
@@ 4 When to Trust Fine-tuned Judges?
(a) JudgeLM-7B applied on SALAD-Bench
(b) Auto-J-13B applied on SALAD-Bench
Figure 4: Accuracy of judge models when applied to different buckets of data grouped with reliability scores.
To verify the effectiveness of the reliability score, we re-conduct the cross-validation in Section 3, wherein the test set was split into halves based on different indicators. As shown in Table 9, our proposed SoftEnt-Cali managed to select the samples with better reliability, therefore achieves a higher accuracy than random split. On the other hand, split based solely on the SoftEnt or perplexity of the judge model underperforms, as they did not exclude the influence of the foundation model.
We conducted an additional experiment wherein samples were grouped into buckets based on their SoftEnt-Cali scores, allowing us to assess the judge model across different buckets. As shown in Figure 4, the accuracy of judge models exhibit a strong correlation with the SoftEnt-Cali score. Notably, for both judges on SALAD-bench, buckets with higher scores demonstrated an accuracy of 90%, whereas those with lower scores performed akin to random guessing. This underscores the efficacy of SoftEnt-Cali as a reliability indicator of the judge model’s performance for evaluation.
The reliability indicator serves as a useful tool to compensate for the limitations and improve the LLM evaluation. For example, it can be used to decide whether an evaluation sample requires human intervention, or to select the most reliable one when multiple judge models are available.
@@ 5 Conclusion
In this work, we conduct an empirical study on the judge models for LLM evaluation. As revealed in our experiments, despite achieving superior evaluation performance on the in-domain testset, the fine-tuned judge models underperforms GPT-4 in terms of several aspects by a large margin, which we believes originates from the task-specific fine-tuning process. We also propose an indicator to quantify the reliability of fine-tuned judges.
Although it is possible to incorporate more diverse fine-tuning data to amend the limitations, as the potential of LLM extends beyond boundaries, there would always be new domains and tasks not covered by the fine-tuning process. Therefore, to draw a conclusion, the fine-tuned judge model cannot serve as a general substitution for GPT-4 in terms of LLM evaluation. It is advisable to exercise caution when leveraging fine-tuned judge models for evaluation in real applications, watching for the overlap between the evaluation scenario and the fine-tuning process.
@@
Our work still has some limitations: 1) The reliability score proposed in our paper can be combined with human inspection or model ensemble, to further improve the LLM evaluation pipeline. For example, when applying judge model for LLM evaluation, we can select the least reliable samples to be evaluated by human evaluator, therefore improving the evaluation accuracy with minimal expense. Due to time and resource constraints , we have to leave this as future work. 2) The work of Zeng et al. (2023) is only a general assessment of evaluator bias, and we did not include fine-grained assessment for different biases, such as position bias Wang et al. (2023a), verbosity bias Saito et al. (2023), etc. 3) Due to time constraints, we did not incorporate manual inspection into the meta-evaluation process. Including human evaluators would enhance the credibility of our claims.
@@ A.1 Training Settings
As mentioned in Section 3, we fine-tune our own judge models based on the four groups of data (JudgeLM Zhu et al. (2023b), PandaLM Wang et al. (2024), Auto-J Li et al. (2023a), Prometheus Kim et al. (2023)), both in generation-style and in classification-style, for the purpose of comparison.
We train all the models on NVIDIA A100-80GB GPUs with Huggingface-transformers Wolf et al. (2020) and DeepSpeed Rasley et al. (2020). Detailed hyper-parameters are presented in Table 10. Notice when comparing generation and classification models, we adopt the same prompt template and same hyper-parameters, with the only difference lies in the prediction method, as illstrated in Figure 5. For generation model, the prediction head reused the pretrained language model head, and is trained akin to the process of language modeling. For classification (regression) model, the prediction head is newly initialized as a linear projection layer, and is decoupled from the language modeling process [^9] Please refer to the class AutoModelForSequence Classification in Huggingface library for more details..
@@ A.2 Prompt Templates
    For Section 3.2, we adopt the prompts presented in Figure 6 to 13 for cross validation. Notice for JudgeLM and PandaLM, their predefined prompts are in the form of pairwise selection, and we make slight modifications to apply them on pointwise grading. For Prometheus, the predefined prompt is in the form of pointwise grading, and we make slight modifications to apply it on pairwise selection. For Auto-J, they predefined prompts both for pairwise selection and pointwise grading. We also adopt the prompts presented from Figure 14 to 17 on MT-Bench, which are all adapted to multi-turn evaluation.
@@
    For Section 3.3, we adopt the prompts presented in Figure 6, 8, 10 and 12, as LLMBar is pair-wise selection.
@@
    For Section 3.4, we adopt the prompts presented in Figure 18 to 21 for JudgeLM, PandaLM and Auto-J, respectively. For Prometheus, as its original prompt comprises of scoring rubrics, we simply define the corresponding rubrics for different benchmarks.
@@
    For Section 3.5, we adopt the prompts presented in Figure 22 and Figure 23 for chain-of-thought prompting.

### v3

@@ 3.2.4 No Benefits from Prompt Engineering
    In-context Learning (ICL) Dong et al. (2023): where task demonstrations are integrated into the prompt as illustration. In our work, we randomly select 2-4 demonstrations from the training set based on the max-context length.
@@
    Chain-of-thought (CoT) Wei et al. (2022): where the input prompt is structured in a way that mimics human reasoning. In our work, the judge model is forced to generate the explanation first, then provide a final judgment.
@@
3.3 The Essence of Fine-tuned Judge: A Task-specific Classifier
@@
Model SALAD-Bench JudgeLM-test PandLM-test Auto-J-test Average
accuracy F1 accuracy F1 accuracy F1 agreement
GPT-4-1106 98.75 65.55 84.24 72.83 75.78 71.51 56.90 78.92
JudgeLM-7B 82.81 57.46 82.39 72.97 68.17 65.18 45.33 69.68
+CascadedEval 96.35 64.56 84.53 71.69 75.08 71.19 54.09 77.51
Auto-J-13B 72.76 52.72 77.79 62.64 72.17 64.10 53.59 69.08
+CascadedEval 93.85 62.73 81.07 64.91 76.27 72.09 57.69 77.22
Table 8: Experiment results of our proposed CascadedEval. We use SE-Cali to quantify the confidence of fine-tuned judge models, and then allocate the less confident 50% to be evaluated by GPT-4.
Model Indicator SALAD-Bench JudgeLM-test PandLM-test Auto-J-test Average
accuracy F1 accuracy F1 accuracy F1 agreement
JudgeLM-7B random 82.50 57.36 80.40 71.87 67.54 65.20 43.25 59.42
perplexity 88.09 59.57 85.48 74.38 72.34 63.25 49.57 61.69
SE 95.10 63.62 89.40 78.09 77.96 64.01 55.89 65.40
SE-Cali 91.56 61.47 88.70 76.30 79.76 67.05 56.32 65.29
Auto-J-13B random 72.29 48.10 77.27 60.7 71.54 63.51 46.12 54.61
perplexity 77.47 52.56 80.28 61.85 75.55 65.84 50.35 57.65
SE 81.98 54.68 79.13 63.58 80.16 66.61 53.01 59.47
SE-Cali 82.08 54.75 80.19 63.58 81.36 70.00 52.30 60.16
Table 9: Comparison of different confidence indicators for the judge models. We split the test sets into halves based on different indicators, and report the performance of the judge on the half with higher scores.
Figure 4: The overall framework of CascadedEval. Conf. denotes the confidence score qualified by our indicator.
@@
Although prior research on instruction-tuning all emphasizes the importance of data diversity Zhou et al. (2023); Lu et al. (2024), the fine-tuning of judges is doing the opposite thing. Therefore, after fine-tuning for a single task with a fixed prompt template, the model lost its generalization ability, and degenerate into a task-specific classifier, which exhibits several limitations due to overfitting.
4 Integrate Fine-tuned Judge with GPT-4
4.1 CascadedEval
Despite the limitations revealed in our study, we aim not to disregard the significance of fine-tuned judges entirely. While LLMs are becoming increasingly prevalent, task-specific models are still wildly used in real applications, considering their superior in-domain performance and cost-effectiveness. Therefore, to make the most of fine-tuned judges for LLM evaluation, we propose a novel method, CascadedEval, integrating proprietary models such as GPT-4 to compensate for the limitations.
The framework of CascadedEval is shown in Figure 4. When applying fine-tuned judge for evaluation, CascadedEval derive the confidence scores as a by-product and allocate the less confident samples to be re-evaluated by GPT-4.
The key of our method is the confidence indicator. While previous works resort to perplexity or self-reflection to quantify the confidence Jung et al. (2024), in this work, we propose an effective confidence indicator based on softmax probability distribution. Given an instruction $x$ and a fine-tuned judge model with parameters $\theta$, the confidence of evaluation $y$ can be estimated as:
$\textrm{SE}(y|x,\theta)=-\frac{1}{T}\sum_{t=1}^{T}\sum_{v=1}^{V}p_{\theta}(y_{t}^{v})\mathrm{log}p_{\theta}(y_{t}^{v}),$ (1)
where $p(y_{t})$ represents the conditional distribution $p(y_{t}|x,y_{<t},\theta)$ at each decoding step, $T$ is the response length, and $V$ is the vocabulary size. If the majority of the probability mass is concentrated on a limited number of vocabulary words, it indicates that the model is confident and the evaluation is more likely to be accurate, and vice versa.
As we would like to quantify whether the sample lies in the task-specific fine-tuning scope, we further calibrate the SE as follows:
@@ 4.1 CascadedEval
With the effective confidence indicator, CascadedEval enhances the accuracy of fine-tuned judges by routing the samples that fall outside the fine-tuning scope to GPT-4. Conversely, compared to using GPT-4 for all evaluations, CascadedEval significantly reduces the API expense without leading to performance degradation. This is combining the best of both worlds.
4.2 Experiments
In this part, we verify the effectiveness of our proposed CascadedEval on both in-domain and out-of-domain meta-evaluation testsets. As shown in Table 8, by allocating 50% of the less confident samples to GPT-4, CascadedEval achieves an accuracy on par with GPT-4, even surpassing GPT-4 sometimes. This verifies that the fine-tuned judges are highly effective when the evaluation sample aligns with the fine-tuning scope. With our proposed confidence indicator, the degree of this alignment can be accurately quantified, allowing for the targeted processing of less confident cases using enhanced proprietary models.
We also conducted an ablation study, where we control the proportion of samples allocated to GPT-4. As shown in Figure 5, with more GPT-4 involvement, the evaluation accuracy firstly increases and then slightly decreases. This notifies that there exists a balance in CascadedEval between fine-tuned judge and GPT-4, which should be carefully monitored to ensure maximum benefit.
Moreover, to verify the effectiveness of the confidence indicator, we re-conduct the cross-validation in Section 3, wherein the test set was split into halves based on different indicators [^9] We did not compare with self-reflection, as the judge model has failed the general instruction-following ability.. As shown in Table 9, our proposed indicators managed to select the samples with higher accuracy. This verifies the effectiveness of softmax distribution-based indicator for confidence estimation.
In an additional experiment, we group the test set into five distinct buckets based on the SE-Cali scores of each sample, and evaluate the accuracy for each bucket separately. As shown in Figure 6, the accuracy on each bucket exhibit a strong correlation with the SE-Cali score. This underscores the efficacy of SE-Cali as a reliable confidence indicator of the judge model.
Figure 5: The variation of accuracy on PandaLM-test with more samples allocated to GPT-4.
Figure 6: Accuracy of Auto-J on SALAD-Bench when applied to different buckets of data grouped by SE-Cali scores, highest in bucket 1 while lowest in bucket 5.
@@ 5 Conclusion
In this work, we conduct an empirical study on the judge models for LLM evaluation. As revealed in our experiments, despite achieving superior evaluation performance on the in-domain testset, the fine-tuned judge models still present several limitations compared with GPT-4.
As the potential of LLM extends beyond boundaries, there will always be new domains and tasks not covered by the fine-tuning process. Therefore, to draw a conclusion, the fine-tuned judge model cannot serve as a general substitution for GPT-4 in terms of LLM evaluation. It is advisable to exercise caution when leveraging fine-tuned judge models for evaluation in real applications, and introduce proprietary models when necessary.
@@
Our work still has some limitations: 1) The CascadedEval proposed in our paper can be combined with human annotator in a semi-supervised manner, to further improve the LLM evaluation pipeline. For example, we can select the least confident samples to be evaluated by human evaluator, therefore ensuring the evaluation accuracy with minimal expense. Active learning can also be introduced, leading to more targeted fine-tuning on the less confident samples. Due to time and resource constraints, we have to leave this as future work. 2) The work of Zeng et al. (2023) is only a general assessment of evaluator bias, and we did not include fine-grained assessment for different biases, such as position bias Wang et al. (2023a), verbosity bias Saito et al. (2023), etc. 3) Due to time constraints, we did not incorporate manual inspection into the meta-evaluation process. Including human evaluators would enhance the credibility of our claims.
@@ A.1 Training Settings
As mentioned in Section 3, we fine-tune our judge models based on the four groups of data (JudgeLM Zhu et al. (2023b), PandaLM Wang et al. (2024), Auto-J Li et al. (2023a), Prometheus Kim et al. (2023)), both in generation-style and in classification-style, for the purpose of comparison.
We train all the models on NVIDIA A100-80GB GPUs with Huggingface-transformers Wolf et al. (2020) and DeepSpeed Rasley et al. (2020). Detailed hyperparameters are presented in Table 10. Notice when comparing generation and classification models, we adopt the same prompt template and same hyper-parameters, with the only difference lying in the prediction method, as illustrated in Figure 7. For generation model, the prediction head reused the pretrained language model head and is trained akin to the process of language modeling. For classification (regression) model, the prediction head is newly initialized as a linear projection layer, and is decoupled from the language modeling process [^10] Please refer to the class AutoModelForSequence Classification in Huggingface library for more details..
@@ A.2 Prompt Templates
    For Section 3.2.1, we adopt the prompts presented in Figure 8 to 15 for cross validation. Notice for JudgeLM and PandaLM, their predefined prompts are in the form of pairwise selection, and we make slight modifications to apply them on pointwise grading. For Prometheus, the predefined prompt is in the form of pointwise grading, and we make slight modifications to apply it on pairwise selection. For Auto-J, they predefined prompts both for pairwise selection and pointwise grading. We also adopt the prompts presented from Figure 16 to 19 on MT-Bench, which are all adapted to multi-turn evaluation.
@@
    For Section 3.2.2, we adopt the prompts presented in Figure 8, 10, 12 and 14, as LLMBar is pair-wise selection.
@@
    For Section 3.2.3, we adopt the prompts presented in Figure 20 to 23 for JudgeLM, PandaLM and Auto-J, respectively. For Prometheus, as its original prompt comprises of scoring rubrics, we simply define the corresponding rubrics for different benchmarks.
@@
    For Section 3.2.4, we adopt the prompts presented in Figure 24 and Figure 25 for chain-of-thought prompting.